A magnet produces a 0.5-T field between its poles, directed to the east. A dust particle is moving straight down at 0.3 cm/s in this field. What is the charge on this particle, if it is experiencing magnetic force 7.1x10-12 N?

The charge should be given with the correct sign.
Answer #1

Magnetic force

F=q(v X B)

Charge on the particle is

q=(F/VB) =(7.1*10-12)/(0.3*10-2*0.5)

q=4.73*10-9C -or- 4.73 nC